A discrete random variable X has the following probability distribution:
x 1 2 3 4
P(X=x) 1/3 1/3 k 1/4
where k is a constant.
(a) Find the value of k.
(b) Find P(X ≤3).
1
Expert's answer
2019-12-24T14:27:29-0500

(a):

"\sum P(x)=1"

So,

"\frac {1}{3} +\frac {1}{3} +k+\frac {1}{4} =1"

"\frac {4+4+3}{12} +k=1"

"\frac {11}{12} +k=1"

"k=1-\frac {11}{12}"

"k=\frac {1}{12}"

(b):

"P(X\le3)=P(X=1)+P(X=2)+P(X=3)"

"P(X\le3)=\frac{1}{3}+\frac{1}{3}+\frac{1}{12}"

"P(X\le3)=\frac{4+4+1}{12}"

"P(X\le3)=\frac{9}{12}"

"P(X\le3)=\frac{3}{4}"